A motorcycle has a constant acceleration of 2.5 m/s2. Both the velocity and acceleration of the motorcycle point in the same direction. How much time is required for the motorcycle to change its speed from (a) 21 to 31 m/s, and (b) 51 to 61 m/s?
1
Expert's answer
2020-09-03T14:10:17-0400

a)


"t_1=\frac{v_f-v_i}{a}=\frac{31-21}{2.5}=4.0\ s"

b)


"t_2=\frac{u_f-u_i}{a}=\frac{61-51}{2.5}=4.0\ s"